OpenStudy (anonymous):

n^2*n^3 that * means multiply

OpenStudy (agent47):

when u multiply two exponentials with the same base, u just add the exponents, so: n^2 * n^3 = n^(2+3)=n^5

OpenStudy (anonymous):

thanks :)